OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

OpenStudy (shamil98):

|x| - 8 = -5 |x| = 3 +(x) = 3 -(x) = 3 x = 3, x = -3

OpenStudy (anonymous):

x= +3 or x= -3

OpenStudy (nincompoop):

AND

OpenStudy (anonymous):

is this inequality?